# Unifying the Three Algebraic Approaches to the CSP via Minimal Taylor Algebras

### v5

@@ 1 Introduction
The Constraint Satisfaction Problem (CSP) has attracted much attention from researchers in various disciplines. One direction of the CSP research has been greatly motivated by the so-called Dichotomy Conjecture of Feder and Vardi [37, 38] that concerns the computational complexity of CSPs over finite relational structures. The Constraint Satisfaction Problem over a finite relational structure $\mathbb{A}$ of finite signature (also called a template), in its logical formulation, is the problem of deciding the validity of a given primitive positive sentence (pp-sentence), i.e., a sentence that is an existentially quantified conjunction of atomic formulas over $\mathbb{A}$ – the constraints. Examples of problems in this class include satisfiability problems, graph coloring problems, and solving systems of equations over finite algebraic structures (see [41, 44, 9, 52]). The CSP is also ubiquitous in artificial intelligence [36].
A classic result in the field is a theorem by Schaefer [57] that completely classifies the complexity of CSPs over relational structures with a two-element domain, so-called Boolean structures, by providing a dichotomy theorem: each such a CSP is either solvable in polynomial time or is NP-complete. The Dichotomy Conjecture of Feder and Vardi states that Schaefer’s result extends to arbitrary finite domains. This conjecture inspired a very active research program in the last 20 years, culminating in a positive resolution independently obtained by Bulatov [28] and Zhuk [59, 60]. The exact borderline between tractability and hardness can be formulated as follows [31, 10, 9].
@@
It was already recognized in Schaefer’s work (in fact, it was the basis of his approach) that the complexity of a CSP depends only on the set of relations that are pp-definable (i.e., definable by a primitive positive formula) from the template. Such sets of relations are now usually referred to as relational clones. The impetus of rapid development in the area after Feder and Vardi’s seminal work [38] was a series of papers [44, 45] that brought to attention and applied a Galois connection between operations and relations studied in the sixties [40, 16], which gives a bijective correspondence between relational clones and clones – sets of term operations of algebras.
@@
One way to phrase this core fact is as follows: for any finite algebra $\mathbf{A}$, its set of invariant relations (subuniverses of powers or subpowers in algebraic terminology) is always a relational clone; every relational clone is of this form; and two algebras have the same relational clone of subpowers if and only if they have the same set of term operations (see Subsection 2.1). For instance, a Boolean CSP, say over the domain $\{0,1\}$, is solvable in polynomial time if and only if the relations of the template are subpowers of one of four types of algebras – an algebra with a single constant operation, a semilattice, the majority algebra, or the affine Mal’cev algebra of $\mathbb{Z}/{2}$ (see Subsection 2.2).
@@
This connection between relations and operations allowed researchers to apply techniques from Universal Algebra. Application of these techniques became known as the algebraic approach to the CSP, although one may argue that the name misses the point a little – the success of the approach lies mostly in combining and moving back and forth between the relational and algebraic side, and this is the case for this paper as well. The general theory of the CSP was further refined in subsequent papers [31, 10] and turned out to be an efficient tool in other types of constraint problems including the Quantified CSP [17, 34, 62], the Counting CSP [30, 25], some optimization problems, e.g. the Valued CSP [51] and robust approximability [6], infinite-domain CSPs [15, 14], related promise problems such as “approximate coloring” and the Promise CSP [20, 2], and many others.
@@
One useful technical finding of [31] is that every CSP is equivalent to a CSP over an idempotent template, i.e. a template that contains all the singleton unary relations. This allows us to use parameters in pp-definitions and omit homomorphic equivalence in the first item of Theorem 1.1. On the algebraic side, this allows us to concentrate on so-called idempotent algebras (see Subsection 7.2). Another important contribution of that paper was a conjecture postulating, for idempotent structures, the exact borderline between polynomial solvability and NP-hardness, which coincides with the borderline stated in Theorem 1.1. The hardness part was already dealt with in the same paper and what was left was the tractability part. Within the realm of idempotent structures, the algebras corresponding to the second item of Theorem 1.1 are so-called Taylor algebras (see Subsection 2.3). The following theorem is therefore the core of the two proofs of the Dichotomy Conjecture.
Theorem 1.2 ([28, 59, 60]).
@@
Let $\mathbb{A}$ be an idempotent structure. If there exists an idempotent Taylor algebra $\mathbf{A}$ such that all relations in $\mathbb{A}$ are subpowers of $\mathbf{A}$, then the CSP over $\mathbb{A}$ is solvable in polynomial time.
@@
Partial results toward Theorem 1.2 include dichotomies for various classes of relational structures and algebras (e.g. the class of 3-element algebras [23] and the class of structures containing all unary relations [24]), understanding of the limits of algorithmic techniques (e.g. local consistency methods [5] and finding generators for the set of solutions [43]), and finding potentially useful characterizations of Taylor algebras (e.g. by means of weak near-unanimity operations [54] and by means of cyclic operations [4]). The papers [5] and [4] initiated a technique which is now referred to as the absorption theory [7]. Absorption theory is one of the fruits of CSP-motivated research which also impacted other CSP-related problems as well as universal algebra (e.g. [1]) and it is one of the three theories this paper is concerned with.
@@
Bulatov and Zhuk in their resolution of the Dichotomy Conjecture (and their prior and subsequent work) developed novel techniques, which we refer to as Bulatov’s theory and Zhuk’s theory in this paper. These theories are (understandably) mostly focused on the task at hand, to prove Theorem 1.2, and as such have several shortcomings. First, some of the new concepts are still evolving as the need arises and they do not yet feel elegant and settled. Moreover, the theories are technically complex which makes it difficult to master them and to apply them in different contexts. This is best witnessed by the absence of results which employ these theories from different authors (needless to say they have already witnessed their potential). Second, they both employ the following trick. Instead of studying a general, possibly wild Taylor algebra, one can first tame it by taking a certain Taylor reduct – an algebra whose operations are only some of the term operations but which is still Taylor. Taking reducts does not result in any loss of generality in Theorem 1.2, since reducts keep all the original invariant relations, so proving tractability for a reduct is sufficient for tractability for the original problem. However, taking reducts does result in loss of generality of the theory and it is not yet clear to which natural classes of algebras the theories apply. Moreover, these reducts are different in the two approaches. Third, connections between Bulatov’s and Zhuk’s theories were not understood at all. While Zhuk’s theory and absorption theory at least had some concepts in common, Bulatov’s theory seemed quite orthogonal to the rest.
@@ 1.1 Taylor algebras
The central concept in absorption theory is that of absorbing subuniverses introduced formally in Subsection 3.2. These are invariant subsets of algebras with an additional property resembling ideals in rings. A fundamental theorem, the absorption theorem [4], shows that nontrivial absorbing subuniverses in Taylor algebras exist under rather mild conditions and this fact makes the theory applicable in many situations. For instance, the strategy in [5] to provide a global solution to a locally consistent instance is to propagate local consistency into proper absorbing subuniverses. The abundance of absorption provided by the absorption theorem makes this propagation often possible, and if it is not, gives us sufficient structural and algebraic information about the instance which makes the propagation possible nevertheless, until the instance becomes trivially solvable.
@@
Zhuk’s starting point is a theorem stating that every Taylor algebra has a proper subuniverse of one of four special types (see Subsections 3.3). Zhuk derives the four types theorem [60] from a complicated result in clone theory, Rosenberg’s classification of maximal clones [56] (the dependence of this approach on Rosenberg’s result is removed in [61]). Given the four types theorem, the overall strategy for the polynomial algorithm for Theorem 1.2 is natural and similar in spirit to the absorption technique – to keep reducing to one of such subuniverses until the problem becomes trivial. Although Zhuk’s theory has a nontrivial intersection with the absorption theory, these connections were not properly explored and verbalized.
@@
Bulatov’s algorithm in his proof of Theorem 1.2 employs a similar general idea, he reduces the instance to certain subuniverses. However, these special subuniverses are defined, as opposed to absorption and Zhuk’s theories, in a very local way. They are sets that are, in a sense, closed under edges (e.g. strongly connected components) of a labeled directed graph whose vertices are the elements of the algebra. Bulatov introduces three basic kinds of edges (see Subsection 3.1), whose presence indicates that the local structure around the adjacent vertices, namely the subuniverse generated by the two vertices, somewhat resembles the three interesting tractable cases in Schaefer’s Boolean dichotomy. What makes this approach work is a fundamental theorem (Theorem 1 [26], see also [22]), the connectivity theorem, which says that the edges sufficiently approximate the algebra in the sense that the directed graph is connected. The proof uses rather technically challenging constructions involving operations in the algebra.
@@
In Section 4 we first describe some of the connections between absorption theory and Zhuk’s theory, and explain simplifications and refinements that were scattered across literature, including a refinement of the absorption theorem that follows from [59, 60]. We also give two new results improving pieces of the two theories. The major novel contribution of Section 4 is Theorem 4.7, a purely relational fact which roughly states that each “interesting” relation that uses all the domain elements in every coordinate pp-defines a binary relation with the same properties or a ternary relation of a very particular shape. Although the proof is elementary and not very long, it enables us to derive both Zhuk’s four types theorem and Bulatov’s connectivity theorems as corollaries. It may be also of interest for some readers to note that theorems in this section often even do not require the algebra to be Taylor – they concern all finite idempotent algebras.
@@ 1.2 Minimal Taylor algebras
The advantage of studying minimal reducts within a class of interest was clearly demonstrated in the work of Brady [18]. He concentrated on so-called bounded width algebras – these are algebras that play the same role in solvability of CSPs by local consistency methods [5] as Taylor algebras do for polynomial time solvability. The theory he developed enabled him to classify all the minimal bounded width algebras on small domains. Our first contributions in Section 5 show that the basic facts for minimal bounded width algebras have their counterparts for minimal Taylor algebras. For instance, Proposition 5.2 shows that every Taylor algebra does have a minimal Taylor reduct, and so minimal Taylor algebras are indeed sufficiently general, e.g., in the CSP context.
@@
Edges, as we already mentioned, are pairs of elements for which the local structure around the pair resembles one of the three interesting polynomially solvable cases in Schaefer’s Boolean dichotomy [57]. More precisely, and specializing to one kind of edges, we say that $(a,b)$ is a majority edge if the subalgebra $\mathbf{E}$ generated by $a$ and $b$ has a proper congruence (i.e., invariant equivalence relation) $\theta$ and a term operation $t$ that acts as the majority operation on the equivalence classes $a/\theta$ and $b/\theta$. The resemblance of the two-element majority algebra is in general quite loose – the equivalence $\theta$ can have many more classes and there may be many more operations in $\mathbf{E}$ other than $t$. However, in minimal Taylor algebras, $\mathbf{E}$ modulo $\theta$ is always term equivalent to the two element majority algebra (Theorem 5.13).
@@ 1.3 Follow up work
Brady has already initiated the project of enumerating minimal Taylor algebras [19]. In particular, he has proved that, up to term-equivalence and permutations of the domain, there are exactly $24$ minimal Taylor algebras on a domain of size $3$. This gives us a concrete list of the hardest tractable CSPs on the 3-element domain, refining the main result of [23]. In an unpublished work of a subset of the authors and Albert Vucaj, the number of cases required to state the 3-element dichotomy was further reduced by means of the pp-constructibility technique from [10].
Another novel contribution in Brady’s CSP notes [19] is the theory of stable sets which provide a common generalization of three out of the four types in Zhuk’s approach. This theory is applied to simplify and improve results on local consistency [5, 49] and robust approximation [6] of CSPs.
@@ 1.4 Organization of the paper
PART I: Main results
@@ 2 Basics
In this section we expand on the explanation of the algebraic approach given in the previous section. The aim is to provide a short but self-contained introduction to the subject which, however, cannot replace more comprehensive material such as [9]. Subsection 2.1 explains the translation of the CSP to the algebraic language, Subsection 2.2 details Schaefer’s dichotomy theorem, and Subsection 2.3 introduces the central concept of the algebraic theory of the CSP, the Taylor algebra. Throughout the paper, notations and concepts are introduced as the need arises, occasionally in a somewhat informal way; please see Section 7 for a more systematic list of definitions.
@@ 2.1 From structures to algebras
Next, we define $\mathscr{C}$ as the set of all operations, i.e., mappings $f:A^{n}\to A$ for some $n$, that preserve all relations in $\mathbb{A}$ (equivalently, all relations in $\mathscr{R}$), i.e., $f$ applied coordinate-wise to tuples in a relation of $\mathbb{A}$ gives a tuple which is again in that relation. Such operations are also called compatible with $\mathbb{A}$ or polymorphisms of $\mathbb{A}$, and we also say that relations in $\mathbb{A}$ are invariant under $f$. The set $\mathscr{C}$ is a clone, that is, it contains all the projections $\operatorname{proj}^{n}_{i}$ (the $n$-ary projection to the $i$-th coordinate) and is closed under composition. Since each $f\in\mathscr{C}$ in particular preserves the singleton unary relations in $\mathbb{A}$, it is idempotent, i.e., $f(a,\dots,a)=a$ for any $a\in A$. Importantly [40, 16], a relation is in $\mathscr{R}$ if and only if it is invariant under every $f\in\mathscr{C}$, so our clone still captures the complexity of the CSP over $\mathbb{A}$.
@@
The final step from clones to algebras is not essential and is taken mostly for convenience: algebras are more classical objects than clones, with a better established terminology. In particular, we can apply the standard constructions of taking subalgebras or subuniverses (universes of subalgebras), products, and quotients over congruences (invariant equivalence relations). This final step is done by selecting some “generating operations” in $\mathscr{C}$ and giving them names. Formally, we take an algebra $\mathbf{A}$ (with universe $A$) of some signature such that the smallest clone containing the operations in $\mathbf{A}$, called the clone of $\mathbf{A}$ and denoted $\operatorname{Clo}(\mathbf{A})$, is equal to $\mathscr{C}$. Note that $\operatorname{Clo}(\mathbf{A})$ is precisely the set of term operations of $\mathbf{A}$, i.e., operations that can be defined from the operations in $\mathbf{A}$ by a term. The algebra $\mathbf{A}$ is not uniquely determined by $\mathbb{A}$, any algebra $\mathbf{A}^{\prime}$ which is term-equivalent to $\mathbf{A}$, i.e., $\operatorname{Clo}(\mathbf{A}^{\prime})=\operatorname{Clo}(\mathbf{A})$, can be taken instead. Also observe that $\mathbf{A}$ is idempotent (consists of idempotent operations) and that the connection to $\mathscr{R}$ remains valid: a relation $R$ is in $\mathscr{R}$ if and only if it is invariant under every operation of $\mathbf{A}$; in other words, if $R$ is a subpower (subuniverse of a finite power) of $\mathbf{A}$. We use $\leq$ to denote the subuniverse or subalgebra relation, e.g., $R\leq\mathbf{A}^{n}$ means that $R$ is a subuniverse of the $n$-th power of $\mathbf{A}$. The set of all subpowers of $\mathbf{A}$ is denoted $\operatorname{Inv}(\mathbf{A})$.
@@
In summary, to every idempotent relational structure $\mathbb{A}$ we assigned an idempotent algebra $\mathbf{A}$ so that the relational clone of all pp-definable relations is equal to $\operatorname{Inv}(\mathbf{A})$, the set of subpowers of $\mathbf{A}$; and the clone of all compatible operations is equal to $\operatorname{Clo}(\mathbf{A})$, the set of all term operations of $\mathbf{A}$. The CSP over $\mathbb{A}$ is polynomial-time equivalent to the CSP over any sufficiently rich structure whose relations are in $\operatorname{Inv}(\mathbf{A})$. It also follows that the CSP over some structure $\mathbb{B}$ is polynomial-time reducible to the CSP over $\mathbb{A}$ whenever $\mathbf{A}$ is a reduct of an algebra $\mathbf{B}$ assigned to $\mathbb{B}$, i.e., $\operatorname{Clo}(\mathbf{A})\subseteq\operatorname{Clo}(\mathbf{B})$.

### v6

@@ 1 Introduction
The Constraint Satisfaction Problem (CSP) has attracted much attention from researchers in various disciplines. One direction of the CSP research has been greatly motivated by the so-called Dichotomy Conjecture of Feder and Vardi [Feder93:monotone, Feder98:monotone] that concerns the computational complexity of CSPs over finite relational structures. The Constraint Satisfaction Problem over a finite relational structure $\mathbb{A}$ of finite signature (also called a template), in its logical formulation, is the problem of deciding the validity of a given primitive positive sentence (pp-sentence), i.e., a sentence that is an existentially quantified conjunction of atomic formulas over $\mathbb{A}$ – the constraints. Examples of problems in this class include satisfiability problems, graph coloring problems, and solving systems of equations over finite algebraic structures (see [Hell04:homomorphism, Jeavons97:closure, Barto17:polymorphisms, Larose06:taylor]). The CSP is also ubiquitous in artificial intelligence [Dechter03:processing].
A classic result in the field is a theorem by Schaefer [Schaefer] that completely classifies the complexity of CSPs over relational structures with a two-element domain, so-called Boolean structures, by providing a dichotomy theorem: each such a CSP is either solvable in polynomial time or is NP-complete. The Dichotomy Conjecture of Feder and Vardi states that Schaefer’s result extends to arbitrary finite domains. This conjecture inspired a very active research program in the last 20 years, culminating in a positive resolution independently obtained by Bulatov [Bulatov17:dichotomy] and Zhuk [Zhuk17:proof, Zhuk20:dichotomy]. The exact borderline between tractability and hardness can be formulated as follows [Bulatov05:classifying, Barto18:wonderland, Barto17:polymorphisms].
@@
It was already recognized in Schaefer’s work (in fact, it was the basis of his approach) that the complexity of a CSP depends only on the set of relations that are pp-definable (i.e., definable by a primitive positive formula) from the template. Such sets of relations are now usually referred to as relational clones. The impetus of rapid development in the area after Feder and Vardi’s seminal work [Feder98:monotone] was a series of papers [Jeavons97:closure, Jeavons98:algebraic] that brought to attention and applied a Galois connection between operations and relations studied in the sixties [Geiger68:pol-inv, Bodnarcuk69:pol-inv], which gives a bijective correspondence between relational clones and clones – sets of term operations of algebras.
@@
One way to phrase this core fact is as follows: for any finite algebra , its set of invariant relations (subuniverses of powers or subpowers in algebraic terminology) is always a relational clone; every relational clone is of this form; and two algebras have the same relational clone of subpowers if and only if they have the same set of term operations (see Subsection 2.1). For instance, a Boolean CSP, say over the domain $\{0,1\}$, is solvable in polynomial time if and only if the relations of the template are subpowers of one of four types of algebras – an algebra with a single constant operation, a semilattice, the majority algebra, or the affine Mal’cev algebra of $\mathbb{Z}/{2}$ (see Subsection 2.2).
@@
This connection between relations and operations allowed researchers to apply techniques from Universal Algebra. Application of these techniques became known as the algebraic approach to the CSP, although one may argue that the name misses the point a little – the success of the approach lies mostly in combining and moving back and forth between the relational and algebraic side, and this is the case for this paper as well. The general theory of the CSP was further refined in subsequent papers [Bulatov05:classifying, Barto18:wonderland] and turned out to be an efficient tool in other types of constraint problems including the Quantified CSP [Borner09:complexity, Carvalho17:complexity, Zhuk22:qcsp], the Counting CSP [Bulatov07:towards, Bulatov13:complexity], some optimization problems, e.g. the Valued CSP [Krokhin17:complexity] and robust approximability [Barto16:robustly], infinite-domain CSPs [Bodirsky18:discrete-temporal, bodirskyBook], related promise problems such as “approximate coloring” and the Promise CSP [Brakensiek18:promise, Bulin19:algebraic], and many others.
@@
One useful technical finding of [Bulatov05:classifying] is that every CSP is equivalent to a CSP over an idempotent template, i.e. a template that contains all the singleton unary relations. This allows us to use parameters in pp-definitions and omit homomorphic equivalence in the first item of Theorem 1.1. On the algebraic side, this allows us to concentrate on so-called idempotent algebras (see Subsection 7.2). Another important contribution of that paper was a conjecture postulating, for idempotent structures, the exact borderline between polynomial solvability and NP-hardness, which coincides with the borderline stated in Theorem 1.1. The hardness part was already dealt with in the same paper and what was left was the tractability part. Within the realm of idempotent structures, the algebras corresponding to the second item of Theorem 1.1 are so-called Taylor algebras (see Subsection 2.3). The following theorem is therefore the core of the two proofs of the Dichotomy Conjecture.
Theorem 1.2 ([Bulatov17:dichotomy, Zhuk17:proof, Zhuk20:dichotomy]).
@@
Let $\mathbb{A}$ be an idempotent structure. If there exists an idempotent Taylor algebra such that all relations in $\mathbb{A}$ are subpowers of , then the CSP over $\mathbb{A}$ is solvable in polynomial time.
@@
Partial results toward Theorem 1.2 include dichotomies for various classes of relational structures and algebras (e.g. the class of 3-element algebras [Bulatov06:three-element] and the class of structures containing all unary relations [Bulatov11:conservative]), understanding of the limits of algorithmic techniques (e.g. local consistency methods [Barto14:local] and finding generators for the set of solutions [Idziak10:few]), and finding potentially useful characterizations of Taylor algebras (e.g. by means of weak near-unanimity operations [Maroti08:weak-nu] and by means of cyclic operations [Barto12:cyclic]). The papers [Barto14:local] and [Barto12:cyclic] initiated a technique which is now referred to as the absorption theory [Barto17:absorption]. Absorption theory is one of the fruits of CSP-motivated research which also impacted other CSP-related problems as well as universal algebra (e.g. [Barto18:CM]) and it is one of the three theories this paper is concerned with.
@@
Bulatov and Zhuk in their resolution of the Dichotomy Conjecture (and their prior and subsequent work) developed novel techniques, which we refer to as Bulatov’s theory and Zhuk’s theory in this paper. These theories are (understandably) mostly focused on the task at hand, to prove Theorem 1.2, and as such have several shortcomings. First, some of the new concepts are still evolving as the need arises, and they do not yet feel elegant and settled. Moreover, the theories are technically complex which makes it difficult to master them and to apply them in different contexts. This is best witnessed by the absence of results which employ these theories from different authors (needless to say they have already witnessed their potential). Second, they both employ the following trick. Instead of studying a general, possibly wild Taylor algebra, one can first tame it by taking a certain Taylor reduct – an algebra whose operations are only some of the term operations but which is still Taylor. Taking reducts does not result in any loss of generality in Theorem 1.2, since reducts keep all the original invariant relations, so proving tractability for a reduct is sufficient for tractability for the original problem. However, taking reducts does result in loss of generality of the theory, and it is not yet clear to which natural classes of algebras the theories apply. Moreover, these reducts are different in the two approaches. Third, connections between Bulatov’s and Zhuk’s theories were not understood at all. While Zhuk’s theory and absorption theory at least had some concepts in common, Bulatov’s theory seemed quite orthogonal to the rest.
@@ 1.1 Taylor algebras
The central concept in absorption theory is that of absorbing subuniverses introduced formally in Subsection 3.2. These are invariant subsets of algebras with an additional property resembling ideals in rings. A fundamental theorem, the absorption theorem [Barto12:cyclic], shows that nontrivial absorbing subuniverses in Taylor algebras exist under rather mild conditions and this fact makes the theory applicable in many situations. For instance, the strategy in [Barto14:local] to provide a global solution to a locally consistent instance is to propagate local consistency into proper absorbing subuniverses. The abundance of absorption provided by the absorption theorem makes this propagation often possible, and if it is not, gives us sufficient structural and algebraic information about the instance which makes the propagation possible nevertheless, until the instance becomes trivially solvable.
@@
Zhuk’s starting point is a theorem stating that every Taylor algebra has a proper subuniverse of one of four special types (see Subsections 3.3). Zhuk derives the four types theorem [Zhuk20:dichotomy] from a complicated result in clone theory, Rosenberg’s classification of maximal clones [Rosenberg70] (the dependence of this approach on Rosenberg’s result is removed in [zhuk21:strong-subalgs]). Given the four types theorem, the overall strategy for the polynomial algorithm for Theorem 1.2 is natural and similar in spirit to the absorption technique – to keep reducing to one of such subuniverses until the problem becomes trivial. Although Zhuk’s theory has a nontrivial intersection with the absorption theory, these connections were not properly explored and verbalized.
@@
Bulatov’s algorithm in his proof of Theorem 1.2 employs a similar general idea, he reduces the instance to certain subuniverses. However, these special subuniverses are defined, as opposed to absorption and Zhuk’s theories, in a very local way. They are sets that are, in a sense, closed under edges (e.g. strongly connected components) of a labeled directed graph whose vertices are the elements of the algebra. Bulatov introduces three basic kinds of edges (see Subsection 3.1), whose presence indicates that the local structure around the adjacent vertices, namely the subuniverse generated by the two vertices, somewhat resembles the three interesting tractable cases in Schaefer’s Boolean dichotomy. What makes this approach work is a fundamental theorem (Theorem 1 [Bulatov16:graph], see also [Bulatov04:graph]), the connectivity theorem, which says that the edges sufficiently approximate the algebra in the sense that the directed graph is connected. The proof uses rather technically challenging constructions involving operations in the algebra.
@@
In Section 4 we first describe some of the connections between absorption theory and Zhuk’s theory, and explain simplifications and refinements that were scattered across literature, including a refinement of the absorption theorem that follows from [Zhuk17:proof, Zhuk20:dichotomy]. We also give two new results improving pieces of the two theories. The major novel contribution of Section 4 is Theorem 4.7, a purely relational fact which roughly states that each “interesting” relation that uses all the domain elements in every coordinate pp-defines a binary relation with the same properties or a ternary relation of a very particular shape. Although the proof is elementary and not very long, it enables us to derive both Zhuk’s four types theorem and Bulatov’s connectivity theorems as corollaries. It may be also of interest for some readers to note that theorems in this section often even do not require the algebra to be Taylor – they concern all finite idempotent algebras.
@@ 1.2 Minimal Taylor algebras
The advantage of studying minimal reducts within a class of interest was clearly demonstrated in the work of Brady [Brady20:bw]. He concentrated on so-called bounded width algebras – these are algebras that play the same role in solvability of CSPs by local consistency methods [Barto14:local] as Taylor algebras do for polynomial time solvability. The theory he developed enabled him to classify all the minimal bounded width algebras on small domains. Our first contributions in Section 5 show that the basic facts for minimal bounded width algebras have their counterparts for minimal Taylor algebras. For instance, Proposition 5.2 shows that every Taylor algebra does have a minimal Taylor reduct, and so minimal Taylor algebras are indeed sufficiently general, e.g., in the CSP context.
@@
Edges, as we already mentioned, are pairs of elements for which the local structure around the pair resembles one of the three interesting polynomially solvable cases in Schaefer’s Boolean dichotomy [Schaefer]. More precisely, and specializing to one kind of edges, we say that $(a,b)$ is a majority edge if the subalgebra generated by $a$ and $b$ has a proper congruence (i.e., invariant equivalence relation) $\theta$ and a term operation $t$ that acts as the majority operation on the equivalence classes $a/\theta$ and $b/\theta$. The resemblance of the two-element majority algebra is in general quite loose – the equivalence $\theta$ can have many more classes and there may be many more operations in other than $t$. However, in minimal Taylor algebras, modulo $\theta$ is always term equivalent to the two element majority algebra (Theorem 5.13).
@@ 1.3 Follow up work
Brady has already initiated the project of enumerating minimal Taylor algebras [brady:notes]. In particular, he has proved that, up to term-equivalence and permutations of the domain, there are exactly $24$ minimal Taylor algebras on a domain of size $3$. This gives us a concrete list of the hardest tractable CSPs on the 3-element domain, refining the main result of [Bulatov06:three-element]. In an unpublished work of a subset of the authors and Albert Vucaj, the number of cases required to state the 3-element dichotomy was further reduced by means of the pp-constructibility technique from [Barto18:wonderland].
Another novel contribution in Brady’s CSP notes [brady:notes] is the theory of stable sets which provide a common generalization of three out of the four types in Zhuk’s approach. This theory is applied to simplify and improve results on local consistency [Barto14:local, Kozik16:weak-consistency] and robust approximation [Barto16:robustly] of CSPs.
@@ 1.4 Organization of the paper
Part I Main results
@@ 2 Basics
In this section we expand on the explanation of the algebraic approach given in the previous section. The aim is to provide a short but self-contained introduction to the subject which, however, cannot replace more comprehensive material such as [Barto17:polymorphisms]. Subsection 2.1 explains the translation of the CSP to the algebraic language, Subsection 2.2 details Schaefer’s dichotomy theorem, and Subsection 2.3 introduces the central concept of the algebraic theory of the CSP, the Taylor algebra. Throughout the paper, notations and concepts are introduced as the need arises, occasionally in a somewhat informal way; please see Section 7 for a more systematic list of definitions.
@@ 2.1 From structures to algebras
Next, we define $\mathscr{C}$ as the set of all operations, i.e., mappings $f:A^{n}\to A$ for some $n$, that preserve all relations in $\mathbb{A}$ (equivalently, all relations in $\mathscr{R}$), i.e., $f$ applied coordinate-wise to tuples in a relation of $\mathbb{A}$ gives a tuple which is again in that relation. Such operations are also called compatible with $\mathbb{A}$ or polymorphisms of $\mathbb{A}$, and we also say that relations in $\mathbb{A}$ are invariant under $f$. The set $\mathscr{C}$ is a clone, that is, it contains all the projections $\operatorname{proj}^{n}_{i}$ (the $n$-ary projection to the $i$-th coordinate) and is closed under composition. Since each $f\in\mathscr{C}$ in particular preserves the singleton unary relations in $\mathbb{A}$, it is idempotent, i.e., $f(a,\dots,a)=a$ for any $a\in A$. Importantly [Geiger68:pol-inv, Bodnarcuk69:pol-inv], a relation is in $\mathscr{R}$ if and only if it is invariant under every $f\in\mathscr{C}$, so our clone still captures the complexity of the CSP over $\mathbb{A}$.
@@
The final step from clones to algebras is not essential and is taken mostly for convenience: algebras are more classical objects than clones, with a better established terminology. In particular, we can apply the standard constructions of taking subalgebras or subuniverses (universes of subalgebras), products, and quotients over congruences (invariant equivalence relations). This final step is done by selecting some “generating operations” in $\mathscr{C}$ and giving them names. Formally, we take an algebra (with universe $A$) of some signature such that the smallest clone containing the operations in , called the clone of and denoted $\operatorname{Clo}(\mathbf{A})$, is equal to $\mathscr{C}$. Note that $\operatorname{Clo}(\mathbf{A})$ is precisely the set of term operations of , i.e., operations that can be defined from the operations in by a term. The algebra is not uniquely determined by $\mathbb{A}$, any algebra which is term-equivalent to , i.e., $\operatorname{Clo}($, can be taken instead. Also observe that is idempotent (consists of idempotent operations) and that the connection to $\mathscr{R}$ remains valid: a relation $R$ is in $\mathscr{R}$ if and only if it is invariant under every operation of $\mathbf{A}$; in other words, if $R$ is a subpower (subuniverse of a finite power) of . We use $\leq$ to denote the subuniverse or subalgebra relation, e.g., $R\leq$ means that $R$ is a subuniverse of the $n$-th power of . The set of all subpowers of is denoted $\operatorname{Inv}(\mathbf{A})$.
@@
In summary, to every idempotent relational structure $\mathbb{A}$ we assigned an idempotent algebra so that the relational clone of all pp-definable relations is equal to $\operatorname{Inv}($, the set of subpowers of ; and the clone of all compatible operations is equal to $\operatorname{Clo}($, the set of all term operations of . The CSP over $\mathbb{A}$ is polynomial-time equivalent to the CSP over any sufficiently rich structure whose relations are in $\operatorname{Inv}(\mathbf{A})$. It also follows that the CSP over some structure $\mathbb{B}$ is polynomial-time reducible to the CSP over $\mathbb{A}$ whenever is a reduct of an algebra assigned to $\mathbb{B}$, i.e., $\operatorname{Clo}($.